OpenStudy (anonymous):

Simplify completely x+2/7x^2-28

OpenStudy (mathstudent55):

Do you mean to reduce this fraction? \( \dfrac{x + 2}{7x^2 - 28} \)

OpenStudy (mathstudent55):

If so, start by factoring completely the numerator and denominator. Hint: the denominator can't be factored, so start by factoring the denominator completely.

OpenStudy (anonymous):

x+2/7x^2-28=x+2/7*(x+2)*(x-2)=1/7x-14

OpenStudy (raden):

you are right